Hypothetically, lets assume that the US is running a trade deficit, please demonstrate how the equation S + (M-X) = I + (G-T) will be rearranged to reflect this situation. What is the relationship between the balance of trade and the flow of international capital and how does the rearrangement of the model reflect this? Please explain why the equation will be arranged in the way in which it is now arranged.
Answer #1

Supply of financial capital = Demand of financial capital

S+(M-X) = I+(G-T)

In case the Government is showing a trade deficit.

Then, the above equation can be rewritten as:

Trade deficit= Domestic investment - Private domestic saving - Government savings

(M-X) = I - S - (T-G)

In this, domestic investment is more than domestic savings. And for the domestic savings to exceed domestic investment, there is a need for more foreign capital.

As we know, the balance of trade reflects the difference between the value of country's import and its export for a particular financial year.

In this model,

The trade deficit is reflected through the equation (M-X) which denotes that the country's Imports are higher than its exports. In this situation, The domestic investments will be higher than the country's domestic savings and the government's budget which is in surplus in this model (T-G).

Generally, the trade deficit means the inflow of more foreign financial capital. This deficit can be the result of the economic boom which tends to decrease the trade balance.

However, if there is a trade surplus in the balance of trade, then the flow of foreign capital will decrease and extra financial capital will be invested abroad. The economic recession tend to increase the trade balance.

Thus, this situation aptly describes the relationship between the trade balance and the flow of international capital.

This surplus of foreign capital can be utilized in different ways and that is mostly dependent on the actual case of the economy.

1. This additional can be offset by reduced private savings, where the domestic investment and public savings will be constant.

2. It can also result in higher domestic investment but the public and private savings will remain unchanged.

3. the government can also absorb this as borrowing and the domestic saving and investment will remain the same.

A real-time example of this situation has occurred in the US in the year 1990s.